OpenStudy (anonymous):

at what point do the lines represented by the equations 2x+y+1=0 and 4x+y-3=0 intersect?

OpenStudy (anonymous):

(2, -5)

OpenStudy (radar):

Y=-2x-1 y=-4x+3 -2x-1=-4x+3 2x=4 x=2 y=-3

OpenStudy (anonymous):

I agree with the work Radar shows, just when you substitute x back into your original equation, you should get y=-5 instead of -3.